OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 2y = 4 x – 5y = 11

OpenStudy (anonymous):

x + 2y = 4 This is the first equation. You are asked to solve for x. This simply means that we want x by itself on one side. Let's keep it on the left side. Now what we want to do is get rid of that 2y

OpenStudy (anonymous):

To do that you subtract 2y to both sides and that is you answer.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

x=4?

OpenStudy (anonymous):

Remember that for one thing we do on one side we have to do the same exact thing on to the other side. x + 2y -2y = 4 -2y You only have to simplify that.

OpenStudy (anonymous):

Uhhh x – 5(2y + 4) = 11 x – 5(–2y + 4) = 11 (2y + 4) – 5y = 11 (–2y + 4) – 5y = 11

OpenStudy (anonymous):

Id be C?

OpenStudy (anonymous):

yeah good job

OpenStudy (anonymous):

:D